Describe why in perfect competition the market price of a good becomes the marginal revenue (MR) associated with a given individual firm producing and selling one more unit.
Answer #1

In a perfectly competitive firm, the firm is a price taker which means that it cannot charge any price and must charge the given market price.

Since the price remains fixed, the Marginal revenue added by producing one more unit of good remains the same as the price.

Example- if the price is 10 and the firm produces 5 units then total revenue = 10*5 = 50 and when it produces 6 units, then the total revenue is = 10*6 = 10.

MR = change in total revenue/change in output

= 60-50/6-5

= 10 = equal to the given market price